OpenStudy (anonymous):

what is 12x + 6 - 8x - 4x - 3 + 12 simplifyed

OpenStudy (anonymous):

just combine like terms. (12x-8x-4x)+(6-3+12) 12-8-4=0 + 6-3+12=15 0x+15=15

OpenStudy (anonymous):

15